OpenStudy (anonymous):

Calculate the average rate of Mg + 2HCl -> MgCl2 + H2 when after 30s 10ml of H2 gas is formed To solve this I need to find the concentration of H2 after 30s, wel the amount of substance is already given ( we convert 10ml to moles) , the remaining physical quantity is the volume of SOLUTION. Not just the volume of gas H2! How do I get that? anyone who could help?

OpenStudy (anonymous):

In standard conditions, 22.4L gas= 1mol molecule, I think this may can help u

OpenStudy (anonymous):

with that I can convert 10ml to X moles, but to calculate the concentration after 30s we still need the volume of the total solution which is not 10ml. How do we acquire that total volume of solution?

OpenStudy (anonymous):

@@kc_kennylau i cant understand react..'s words very well, can u translate it for me?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (anonymous):

13xyhuang2 Im bothering you oo much with this :p

OpenStudy (anonymous):

nothing im just looking around, bt you only need to calculate the average rate of the reaction. That's enough 10ml H2= xmol and then x/30s= average rate of the reaction (in H2) If want to calculate the average rate of the reaction after 30s, the conditions is not allowed

OpenStudy (anonymous):

x/30s= average rate of the reaction , that is not correct because X is mol and V is L, so we get mole / L and that is not the unit of reaction rates! The unit of reaction rates is Molar/ L

OpenStudy (frostbite):

Okay we define the rate as \[\Large v=\frac{ d[H _{2}] }{ dt }\] For a observable change we write it as: \[\Large v=\frac{ \Delta [H _{2}] }{ \Delta t }=\frac{ \left[ H _{2} \right]_{t2}-\left[ H _{2} \right]_{t1} }{ t _{2}-t _{1} }\] Assume that hydrogen gas behave as a ideal gas and use the ideal gas equation to calculate the amount of moles the gas occupy assume STP.

OpenStudy (anonymous):

I get n= 4.46 x 10^-4 moles

OpenStudy (anonymous):

now I need the volume of the total solution which includes all the other substances involved in the reaction: "Mg + 2HCl -> MgCl2 + H2" . This is the only way to calculate the concentration. If we take for volume of solution 10ml that would be wrong right? because thats only the volume of the H2 gas and not the volume of the solution. And the definition of concentration is moles of specific substance divided by volume of solution ( meaning not just the volume of 1 single gas)

OpenStudy (frostbite):

Yes but you can also give the answer as mol/min.

OpenStudy (frostbite):

It of cuase depend on the one you to turn the assignment into, but yes your considerations are correct.

OpenStudy (anonymous):

Can you tell me how to calculate the volume of the entire system ( Hcl, MgCl2, solid Mg & H2 gas)?

OpenStudy (frostbite):

Hmmm I don't think so sorry.

OpenStudy (anonymous):

oh I just found the answer lol thanks for your help Frostbite!

OpenStudy (frostbite):

No problem, no idea if it was any help what I wrote, but glad you found the answer! :)